OpenStudy (anonymous):

the sum of two consecutive odd integers is -116 what are the integers

OpenStudy (anonymous):

-57 + -59 = -116

OpenStudy (anonymous):

integers are -57 and -59

OpenStudy (xguardians):

x + (x+2) = -116 2x + 2 = -116 2x = -118 x = -59 -59 and -57

OpenStudy (anonymous):

lol thanks @ruben12345

OpenStudy (anonymous):

your welcome :) ii wish i was good in algebra :(